OpenStudy (anonymous):

Hellppp pleaseee /.\ Simplify the following expression: x - 4y + 2z - 3x - 3y + 2z = ?

jimthompson5910 (jim_thompson5910):

combine like terms x-3x = -2x -4y - 3y = -7y 2z + 2z = 4z So that's why x - 4y + 2z - 3x - 3y + 2z simplifies to -2x - 7y + 4z

OpenStudy (iamtheone):

Hello, welcome to Openstudy..... x-4y+2z-3x-3y+2z=(x-3x-3y-4y+2z+2z)=-2x-7y+4z.....

OpenStudy (iamtheone):

Happy Studying!!!!!